OpenStudy (anonymous):

Can someone Please help me a medal will be given! The equation of line EF is y = 1 over 2x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, −2). y = −2x − 2 y = negative 1 over 2x + 2 y = 1 over 2x − 2 y = − 2x + 2

OpenStudy (anonymous):

@TuringTest @whpalmer4 @e.mccormick @hoblos

OpenStudy (whpalmer4):

EF is \[y=\frac{1}{2}x+6\]?

OpenStudy (whpalmer4):

I'm going to assume it is. That is a line equation in slope-intercept form, or \[y = mx+b\]where \(m\) is slope and \(b\) is the y-intercept. You need a parallel line. Parallel lines have the same slope. Finally, you need to write a parallel line that goes through the point (0,-2). To do that, use the point-slope formula for a line with slope \(m\) going through point \((x_1,y_1)\): \[y-y_1 = m(x-x_1)\] After you've written that, solve it for \(y\) and compare with your answer choices.

OpenStudy (anonymous):

can you elaborate a little more @whpalmer4

OpenStudy (whpalmer4):

Ask a specific question. I told you the entire procedure...

OpenStudy (whpalmer4):

Do you understand how to get the slope of the current line?

OpenStudy (anonymous):

no and how do i find y-y1

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

You still haven't answered my original question about whether I interpreted your description of the equation correctly.

OpenStudy (anonymous):

no @whpalmer4

OpenStudy (whpalmer4):

If that isn't the correct equation, what is? Either use the equation editor button or draw it

OpenStudy (anonymous):

wait no you are right @whpalmer4

OpenStudy (whpalmer4):

Okay, thank you. Compare the equation from the problem with the one beneath it: \[y = \frac{1}{2}x+6\]\[y=mx+b\] If they are equal, what are the values of \(m\) and \(b\)?

OpenStudy (anonymous):

@whpalmer4 i really dont understand this

OpenStudy (whpalmer4):

You have two equations right over each other. You can't figure out which numbers and which variables correspond?

OpenStudy (whpalmer4):



OpenStudy (whpalmer4):

 \[b = 6\]Can you figure out what \(m\) is?

OpenStudy (anonymous):

1/2?

OpenStudy (whpalmer4):

Very good! That wasn't so hard, was it?

OpenStudy (whpalmer4):

There are a number of things where you'll do the same operation (comparing two equations, one with numbers and the other with variables, and using the similarity to find the values of the variables) so it's best that you get comfortable with the notion.

OpenStudy (whpalmer4):

So the slope of the line is \(m = \frac{1}2\) As I said earlier, the point-slope formula will allow us to write the equation of a line with slope \(m\) passing through a point \((x_1,y_1)\). We know that \(m = \frac{1}{2}\), and our known point that the line is to pass through is \((x_1,y_1) = (0,-2)\) Any question about that?

OpenStudy (whpalmer4):

Okay, so you know \(m, x_1, y_1\), plug them into the formula: \[y-y_1 = m(x-x_1)\]What do you get?

OpenStudy (anonymous):

so how do i find y-y1

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Take out your pencil. Write "y - " then write in the value of \(y_1\)

OpenStudy (anonymous):

@whpalmer4 and y is 2x?

OpenStudy (whpalmer4):

Geez, you're really not getting this. \[m = \frac{1}{2}\]\[(x_1,y_1) = (0,-2)\]\[x_1 = 0\]\[y_1 = -2\] \[y-y_1 = m(x-x_1)\]\[y - (-2) = \frac{1}{2}(x-0)\]\[y+2 = \frac{1}{2}x\]Can you solve that equation for \(y\) on the left, and everything else on the right?

OpenStudy (anonymous):

i got y = -1/2 over 2x + 2

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Can you write out all of the steps you took to produce that result, starting with my equation?

OpenStudy (anonymous):

after all i put i got y+2=1/2 the i subtracted two from both sides

OpenStudy (whpalmer4):

What happened to the x?

OpenStudy (whpalmer4):

\[y + 2 = \frac{1}{2}x\]Subtract 2 from both sides: \[y+2-2 = \frac{1}{2}x - 2\]\[y = \frac{1}{2}x-2\]